### v2

# These Are Not All the Features You Are Looking For: A Fundamental Bottleneck in Supervised Pretraining

Abstract
Transfer learning is a cornerstone of modern machine learning, promising a way to adapt models pretrained on a broad mix of data to new tasks with minimal new data. However, a significant challenge remains in ensuring that transferred features are sufficient to handle unseen datasets, amplified by the difficulty of quantifying whether two tasks are "related". To address these challenges, we evaluate model transfer from a pretraining mixture to each of its component tasks, assessing whether pretrained features can match the performance of task-specific direct training. We identify a fundamental limitation in deep learning models – an "information saturation bottleneck" – where networks fail to learn new features once they encode similar competing features during training. When restricted to learning only a subset of key features during pretraining, models will permanently lose critical features for transfer and perform inconsistently on data distributions, even components of the training mixture. Empirical evidence from published studies suggests that this phenomenon is pervasive in deep learning architectures – factors such as data distribution or ordering affect the features that current representation learning methods can learn over time. This study suggests that relying solely on large-scale networks may not be as effective as focusing on task-specific training, when available. We propose richer feature representations as a potential solution to better generalize across new datasets and, specifically, present existing methods alongside a novel approach, the initial steps towards addressing this challenge.
1 Introduction
Transfer learning has become a cornerstone of modern machine learning and artificial intelligence (Bottou, 2011; Collobert et al., 2011; Oquab et al., 2014; Bommasani et al., 2021). In its simplest form – supervised transfer learning, a relatively large deep learning network is pre-trained on a large and diverse dataset composed of cheap examples that are somehow related to the task of interest. In a second phase, the network is adapted using a much smaller dataset composed of examples that directly illustrate the task of interest. This second dataset is often too small to train a network of that size from scratch (Li et al., 2017). It is hoped that the representations constructed during the pretraining phase will be useful during the adaptation phase.
It remains an significant challenge to ensure that transferred features are sufficient to handle new, unseen datasets. A pretrained network that is missing crucial feature information may not perform on par with direct learning, that is, a network with all the features we could obtain by training directly on enough data to learn a target task.
Foundation model folklore suggests that, by pretraining these enormous models on "everything", they would learn "all the features (abilities) we could ever want." Should we invest into building ever larger all-purpose foundational models, or into collecting tasks-specific datasets for training smaller specialized models?
Positive findings can only validate transfer from a particular foundational model to a particular target task. We cannot say more because we do not know how to even state that the target task is related to the pre-trained model “abilities”.
Instead, we provide a negative result by exposing shortcomings that appear under very favorable conditions. We assume that the pre-training data is drawn from a mixture of distribution that includes the target task as a component, and we show that we cannot expect that an adapted model will always match a directly trained model.
If we cannot always match direct training when adapting a model pre-trained on a mix that includes the target task, we cannot expect it either when we know even less about the pre-training data.
We uncover an information saturation bottleneck that occurs when models cannot learn new features after they have encoded similar competing features. When networks are restricted to are restricted to learning only a subset of key features during pretraining, they lose crucial feature information that is essential to effectively adapt to new tasks, or even ones they have seen before. These lost features are nearly impossible to recover by simply fine-tuning on limited data.
Although this phenomenon was not previously recognized as such, we argue that it is pervasive in deep learning models. We analyze this phenomenon rigorously through a simple theoretical counterexample, and discuss several published papers whose empirical findings can be explained by a consistent loss of features that would be useful for transfer.
Additionally, factors such as the distribution imbalances or shuffling order of the training can determine which features a model can learn over time. Finally, we propose richer feature representations as a potential solution to improve generalization across new datasets and present existing methods alongside novel approaches which represent the beginning of addressing this challenge.
Section 2 presents a formal framework to assess effectiveness of transfer. Section 3 demonstrates through a simple example, that even in optimal conditions, a deep network might miss key features. Section 4 reveals a fundamental bottleneck of deep networks, supported by empirical evidence across existing studies. It also introduces "rich representations" as a potential solution. Section 5 presents a method of constructing rich representations that improves performance when transferred to unseen distributions, at no additional pretraining cost.
Why study Mixture Distributions?
This paper examines how real-world pretraining works by studying complicated datasets with many imbalanced subgroups. We use a mathematical model called mixture distribution to represent these complex datasets as a single cohesive structure made up of multiple probability distributions.
Mixture of Experts
This is distinct from the neural network architecture known as "mixture of experts" (MoEs), where specialized "expert" models learn different parts of a task. When training on real-world datasets with many components, each "expert" may focus on smaller number tasks, but still be bound by the same limitations.
2 Problem Statement
Consider a set of distributions $\{P^{[j]}(X,Y)\}_{j}$ representing data from different populations, and a mixture distribution $P^{[\mathrm{mix}]}(X,Y)$ formed by a weighted sum of these subdistributions. [^1] Note: This mixture is non-trivial, meaning that all component distributions are present.
$P^{[\mathrm{mix}]}(X,Y)=\sum_{j}\lambda_{j}\cdot P^{[j]}(X,Y)\quad\text{where}~{}\lambda_{j}>0,~{}\sum_{j}\lambda_{j}=1.$
Suppose that we pretrain a network on mixture $P^{[\mathrm{mix}]}$. Let $P^{[i]}$ denote a fixed arbitrary subdistribution. Assuming that we have enough data to directly train a network on $P^{[i]}$, we ask
Question: Do features pretrained using $P^{[\mathrm{mix}]}$ work nearly as well as features learned directly for target $P^{[i]}?$
2.1 Two Ways to Train
Consider a neural network $f$ of the following form
$f(X;\theta,\bm{\gamma})=\sum_{j}\,\underbrace{\varphi_{j}(X;\theta)}_{\text{feature extractor}}\cdot\underbrace{\gamma_{j}}_{\text{classifier}}$ (1)
This network has two main components.
The feature extraction layers ($\varphi_{j}(X;\theta)$) initially transform the raw input $X$ into real-valued features using parameters $\theta.$ Furthermore, these transformations may be non-linear, enabling the extraction of complex signals from the input data.
The linear classification layer ($\gamma_{j}$) combines linearly the extracted features to produce the final prediction. [^2] We explain in the following section why it is sufficient to study the linear case.
There are two ways to train such a model for distribution $P^{[i]}$ (Figure 1)
  1. 1.
    The model could be directly trained on the target distribution $P^{[i]}.$ In this case, the model learns both feature parameters $\theta$ and classifier weights $\bm{\gamma}$ optimized for the target distribution $P^{[i]}$. This yields the final trained model $f(\cdot\ ;\theta^{[i]},\bm{\gamma}^{[i]})$.
  2. 2.
    Alternatively, we could train a model through transfer learning (via linear probing).Footnote 2 The model is pretrained on a mixture distribution $P^{[\mathrm{mix}]}$. The feature parameters $\theta^{[\mathrm{mix}]}$ are frozen and only the linear classifier weights $\bm{\gamma}$ are fine-tuned on $P^{[i]}$. This is called linear probing.
The two approaches primarily differ in how they learn features. In direct training, a model is capable of capturing any representable feature from the target distribution $P^{[i]}$. However, in transfer learning, the model is restricted to selecting features based on a different pretraining task $P^{[i]}$, which includes $P^{[i]}$ as a component among many.
Under the constraints of transfer learning, it is unclear whether models pretrained on broad mixtures retain sufficient features to match the performance task-specific training.
To assess their effectiveness in preserving performance, we explore whether the direct training solutions $f(X;\theta^{[i]},\gamma^{[i]})$ of any component $P^{[i]}$ can be matched (or approximated) by linearly combining features $\varphi_{j}(X;\theta^{[\mathrm{mix}]})$ from pretraining on the mixture.
We assume access to unlimited training data, which is important for effective direct training. Theoretically, we ask whether the function space spanned by features $\varphi_{j}(X;\theta^{[\mathrm{mix}]})$ contains the direct training solutions $f(\dots,\theta^{[i]},\gamma^{[i]})$ trained on unlimited data.
Figure 1: Two ways to train a classifier for $P^{[i]}$: directly from $P^{[i]}$ or transferred from $P^{[\mathrm{mix}]}$.
2.1.1 Studying the Linear Case is Sufficient
Linear Probing:
We focus on the scenario of transfer learning via linear probing, where only the final classification layer is modified during fine-tuning. Thus, our results directly apply to linear probing.
General Fine-Tuning:
Our findings extend beyond linear probing to more common case where all network layers are adjusted. This can be written as a particular instance within our framework.
Consider the general case of fine-tuning a pre-trained neural network $f(X;\mathbf{w}^{[\mathrm{mix}]})$ that updates all of its weights for target $P^{[i]}$. Under the assumptions of fine-tuning, the final weights $\mathbf{w}^{[i]}$ remain close to the pre-trained weights $\mathbf{w}^{[\mathrm{mix}]}.$ Thus, $\mathbf{w}^{[i]}$ can be reasonably approximated as a first-order Taylor expansion around $\mathbf{w}^{[\mathrm{mix}]}$
[…]
Observe that $f(X;\mathbf{w}^{[i]})$ is a linear combination of ${\partial f}/{\partial w_{j}}(X;\mathbf{w}^{[\mathrm{mix}]})$, which are exactly the neural tangent kernel (NTK) features (Jacot et al., 2018). The fine-tuning operation effectively modifies only the linear coefficients $(w_{j}^{[i]}-w_{j}^{[\mathrm{mix}]})$. Thus, this is an instance of our abstract problem statement, in the space of these NTK features.
Non-Linear Classifier:
It is also common to fine-tune multiple non-linear classification layers of a network. The previous argument also applies when we modify some, not all, weights
2.1.2 Networks with Non-Zero Training Error
When training neural networks, it is common practice to optimize the model until the training error reaches zero. However, we cannot learn useful information from studying these.
Transfer has historically been used when there is insufficient data to directly learn a task. A network may achieve zero training error without generalizing perfectly to the entire distribution. When not all features have been learned, transfer to a subdistribution indicates how a network may generalize.
Our problem setup makes the assumption of unlimited training data. When a model has trained on everything, zero training error translates to zero test error. When a network already performs perfectly on all subdistributions, transfer to a component tells us nothing about how it may generalize.
2.1.3 Covariance
Proposition 2.1.
Suppose we find a feature $\varphi(X)$ correlated output $Y$ under the subdistribution $P^{[i]}$ Then, $\varphi(X)$ is correlated with $Y$ under any mixture $P^{[\mathrm{mix}]}=\sum_{j}\lambda_{j}P^{[j]}$, except maybe for a measure zero set of mixture coefficients $(\lambda_{j})$.
$\begin{split}\forall\neq 0\quad\implies\mathrm{Cor}_{P^{[\mathrm{mix}]};\lambda}(\varphi(X),Y)\neq 0\quad\text{for }\text{ a.e. }\mathbf{\lambda}\end{split}$
Proof.
Assume that this is not the case. Two variables are correlated if and only if they have nonzero covariance. Therefore, mixture coefficients $\lambda_{i}$ must satisfy the equation:
$0=\mathrm{Cov}^{[\mathrm{mix}]}(Y,\varphi(X))=\sum_{j}\lambda_{j}\mathrm{Cov}^{[j]}(Y,\varphi(X)).$
Under our assumption, at least one of the covariances $\text{Cov}^{[j]}$ is nonzero. Thus, this equation represents a measure zero subset of mixture coefficients. Outside of this subset, $\mathrm{Cov}^{[\mathrm{mix}]}(Y,\varphi(X))\neq 0$. ∎
Intuitively, the remaining components must exactly cancel out the positive covariance of $P^{[i]}$, which only occurs for precisely-balanced mixture coefficients.
This suggests that a feature interesting to a target $P^{[i]}$ (i.e. - correlated with Y) is also interesting to the pretraining distribution $P^{[\mathrm{mix}]}$. However, a nonzero covariance alone does not mean that this feature will be part of the solution $f(\theta^{[\mathrm{mix}]},\gamma^{[\mathrm{mix}]})$ of the pre-training problem (useful).
If the information $\varphi(X)$ provides about $Y$ has already been introduced by existing features in the solution, it is no longer useful. This is consequential when the training setup is biased towards learning sparse representations.
Sparsity in deep networks
Stochastic gradient learning algorithms in deep learning models often have an implicit sparsity bias (Gunasekar et al., 2017; Andriushchenko et al., 2023), a likely contributor to the effectiveness of deep learning. In an oversimplified account of the analysis of Andriushchenko et al., when using stochastic gradient descent with large step sizes, features are added and removed somewhat randomly, but those that contribute to reducing the training error tend to stick around. Depending on the random order in which features are discovered by the training algorithm, it may construct different sparse solutions.
Many practical regularization tricks such as early-stopping or weight decay have a similar effect. In particular, weight decay acts differently in the last layer of a deep network (spreading the weights over all the available features) than in the inner layers (quickly pruning features that do not rapidly help reducing the training error.)
3 Counterexample
This study explores whether the features learned from mixture distribution perform as well across component tasks $P^{[i]}$ as features learned specifically for each task. We introduce a simple counterexample demonstrating that this desired outcome does not hold.
Example Setup
Imagine a feature extractor which can learn only two feature extraction functions $\Phi(X)=(\varphi_{1}(X),\varphi_{2}(X))$. Depending on parameters $\theta,$ the space spanned by the selected features can have zero, one, or two dimensions.
Consider four individual distributions in this space, illustrated in Figure 3:
  • $P^{[1]}[\>\Phi(X){=}(+1,0),\>Y{=}+1\>]=\tfrac{1}{2}$ and $P^{[1]}[\>\Phi(X){=}(0,\pm 1),\>Y{=}-1\>]=\tfrac{1}{4}$.
  • $P^{[2]}[\>\Phi(X){=}(-1,0),\>Y{=}+1\>]=\tfrac{1}{2}$ and $P^{[2]}[\>\Phi(X){=}(0,\pm 1),\>Y{=}-1\>]=\tfrac{1}{4}$.
  • $P^{[3]}[\>\Phi(X){=}(0,+1),\>Y{=}-1\>]=\tfrac{1}{2}$ and $P^{[3]}[\>\Phi(X){=}(\pm 1,0),\>Y{=}+1\>]=\tfrac{1}{4}$.
  • $P^{[4]}[\>\Phi(X){=}(0,-1),\>Y{=}-1\>]=\tfrac{1}{2}$ and $P^{[4]}[\>\Phi(X){=}(\pm 1,0),\>Y{=}+1\>]=\tfrac{1}{4}$.
Each distribution is defined across three distinct points: two points with probability $1/4$, and the third point with probability $1/2$. Combining these forms a mixture $P^{[\mathrm{mix}]},$ supported over four points (Figure 3).
$P^{[\mathrm{mix}]}=\lambda_{1}\cdot P^{[1]}+\lambda_{2}\cdot P^{[2]}+\lambda_{3}\cdot P^{[3]}+\lambda_{4}\cdot P^{[4]}\quad\text{where}~{}\lambda_{i}>0,~{}\sum_{i=1}\lambda_{i}=1$
A distribution $P^{[\mathrm{mix}]}$ with non-zero mixture coefficients $\lambda_{i}$, therefore, offers positive examples at $\Phi(X)=(\pm 1,0)$ and negative examples at $\Phi(X)=(0,\pm 1)$.
The Optimal Classifier
Suppose that we optimize binary classifiers for each of the four distributions. The training error of each $P^{[i]}$ is minimized using one specific feature: $\varphi_{1}(X)$ for $P^{[1]}$ and $P^{[2]}$, $\varphi_{1}(X)$ for $P^{[3]}$ and $P^{[4]}$.
Now, consider a binary classifier optimized for $P^{[\mathrm{mix}]}.$ Since its four points cannot be linearly separated, the training error is minimized by misclassifying the least-weighted point and correctly classifying the other three. Depending on which point is ignored, the optimal classifier is one of the four solutions in Figure 3
Figure 2: Four subdistributions and a combined mixture distribution are represented by red points (labeled $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+1$) and blue points (labeled $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1$) Each subdistribution includes three points; two points contain an equal number of examples, while the third point contains twice as many. The size of each point reflects the number of examples it represents. The final mixture distribution is a weighted average of these four component distributions (not drawn to scale).
Figure 3: Four cases illustrate optimal classifiers for different mixture distributions. Each classifier uses a dotted line to separate red points (labeled $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+1$) from the blue points (labeled $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1$). Sincef points are not linearly separable, an optimal classifier ignores the least-weighted point (smallest) and classifies the remaining three points. Interestingly, this optimal classifier can be sparsely represented using just one feature, either $\varphi_{1}(X)$ or $\varphi_{2}(X),$ but cannot classify two of the four subdistributions.
Observe that the optimal classifier $P^{[\mathrm{mix}]}$ can be sparsely represented using just one of the two features, determined by the balance of the mixture. Hence, a deep neural network which exhibits sparsity bias will represent this optimal solution by learning just one feature. In other words, the network learns a feature space that does not contain the optimal solutions for two of its four subdistributions. Specifically, when it learns only $\varphi_{1}$, it cannot correctly classify $P^{[3]}$ and $P^{[4]}$; when it learns only $\varphi_{2}$, it cannot correctly classify $P^{[1]}$ and $P^{[2]}$.  [^3] Some mixtures have multiple clusters of equal minimal weight and, therefore, multiple optimal solutions. However, this is a set of negligible probability.
We find that training a deep network (with sparsity bias) on the mixture distribution yields a sparse representation which can classify only half of its subdistributions. Even in this advantageous setup, a model pretrained on a mixture may miss important features for its subtasks.
3.1 Assumptions
In this counterexample, we study the training error as an indicator of how well a model can adapt to its subdistributions. We make the following assumptions
Assumption 1: A Model Will Learn These Features(Papyan et al., 2020) finds that a neural network learns representations which "collapse" into the problem space features, $\varphi_{1}(X)$, $\varphi_{2}(X)$
Assumption 2: We Cannot Study Non-Linear Decision Boundaries A network with a non-linear decision boundary could learn both features and optimally classify all subdistributions. However, we are not interested in studying solutions with zero training error (see Section 2.1.2).
4 True in Practice?
Our simple counterexample illustrates how a model may struggle to learn crucial features during training, which creates biased performance on subtasks of the training distribution.
When an optimally-trained model can miss a crucial feature in such a simple setup, it suggests that deep networks may frequently miss important feature when trained on more complex datasets. In fact, the results of many already-published papers could be explained by this phenomenon. Instead of running redundant experiments, we describe this limitation and explore existing experiments below.
4.1 An Information Saturation Bottleneck
We identify a fundamental limitation of deep learning models that arises from their sparsity bias – an information saturation bottleneck. These networks struggle to learn new features when they have encoded similar competing features during training.
When networks are restricted to learning a subset of key features during pretraining, they permanently lose critical features needed for effective transfer, leading to inconsistent performance on data distributions–even components of the pretraining mixture.
The previous counterexample demonstrates that balance of a mixture may determine which features a deep network will learn over time. We provide empirical evidence for this and identify additional factors which influence the subset of features learned by a network.
4.2 Spurious Features
Table 1: Empirical Evidence (Left) Performance gap after recovering core features with transfer (Right) The simple supervised models trained by their supervised ’DASHA’ workflow outperform eight well Genomic Foundation Models (GFMs) on the diverse NT benchmark (Xu et al., 2025)
Trial Accuracy
Data Direct Transfer via LinProb Change
($99\%$ corr) $P^{[bal]}$ $P^{[\mathrm{mix}]}\to P^{[bal]}$
MNIST-Fashion $97\%$ $94\%$ ($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-3\%$)
MNIST-CIFAR   $90\%$ $81\%$ ($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-9\%$)
Model Model Pretraining Average Average Mean Median
Size Base-Pairs Score $\uparrow$ Rank $\downarrow$ %Imp.$\uparrow$ %Imp.$\uparrow$
Enformer 252M 4B 0.569 11.86 27.73 27.91
NT-1000G (500M) 500M 20.5T 0.625 10.52 33.48 36.74
NT-1000G (2.5B) 2.5B 20.5T 0.656 7.0 36.58 40.86
NT-Multispecies (500M) 500M 174B 0.700 3.81 40.76 45.07
Foundation NT-Multispecies (2.5B) 2.5B 174B 0.697 4.08 40.51 45.52
Models DNABERT-2 117M 32.5B 0.680 6.88 38.65 43.59
HyenaDNA-1K 1.6M 3.2B 0.708 6.92 41.2 43.36
HyenaDNA-32K 1.6M 3.2B 0.630 10.22 33.96 36.93
Caduceus-PS 1.9M 35B 0.689 6.69 39.08 41.38
Caduceus-PH 1.9M 35B 0.725 4.69 42.63 45.01
Supervised Wide ResNet 2.0M 0 0.694 6.83 37.16 43.08
Models UNet 4.5M 0 0.68 7.78 38.67 42.69
DASHA (our workflow) 10.5M 0 0.761 3.69 46.33 49.08
Spurious features are superficial features which are correlated with but do not predict the label. For example, a bird classification model may incorrectly learn to use a water background to predict a waterbird and a land background to predict a land bird.
(Pezeshki et al., 2021) find that learning spurious features hinders the appearance of core features essential for accurate prediction. They find that spurious features to arise from bias towards certain classes in complicated mixtures of data, descriptive of real-world datasets. Core features cannot appear after the gradient is consumed by spurious features, which provide the same information but cannot be used for prediction. Class imbalances produce spurious features which "starve" the model of essential features.
(Kirichenko et al., 2023b) propose a promising mitigation to recover core features via transfer, by reweighting the remaining features without the spurious feature $\varphi_{\text{spu}}$. They consider that a model, pretrained on a biased mixture $P^{[\mathrm{mix}]}$, may still learn core features after learning easier spurious features. They remove $P\varphi_{\text{spu}}$ by transferring the network to a class-balanced subdistribution $P^{[bal]}$ (where $\mathrm{Cor}\big{(}\varphi_{\text{spu}}(X),Y\big{)}=0$). The remaining core features are optimized via linear probing.
Despite significant accuracy improvements, a performance gap $\{3\%,9\%\}$ remains between the adapted model and one trained directly on $P^{[bal]}$ (Table 1). Here, directly training on P[bal] produces a solution which includes all of the core features. In contrast, pre-training on $P^{[\mathrm{mix}]}$ gives a solution which does not. Linear probing on $P^{[bal]}$ cannot recover these features.
In this example of the bottleneck, pretraining on an imbalanced mixture produces misleading spurious features, which prevent the network from learning core features which are essential for prediction.
4.3 Genomic Foundation Models
(Xu et al., 2025) find that simple supervised CNNs can be easily trained to match the performance of transformer-based genomic foundation models (GFMs). Using their "DASHA" workflow, they train and evaluate supervised models on a diverse set of genomic tasks $P^{[\mathrm{adja}]}$ from the Nucleotide Transformer (NT) benchmark. They compare to the performance of well-known GFMs with up to 2.5B parameters, which are pretrained on a broad mixture $P^{[\mathrm{mix}]}$ of genomic tasks. They find the directly trained models to consistently outperform the GFMs across all aggregated metrics of $P^{[\mathrm{adja}]}$ benchmark (Table 1).
Certain GFMs such as Caduceus or NT-Multispecies(500M) perform almost as well as supervised networks on specific tasks of $P^{[\mathrm{adja}]}$, but miss features needed to perform well across all $P^{[\mathrm{adja}]}$. This suggests that large scale models learn some, but not all critical features for $P^{[\mathrm{adja}]}$ which appear in the extensive $P^{[\mathrm{mix}]}.$ While GFMs can excel in specific scenarios, their performance on new tasks is, ultimately, constrained by missing features that arise from information saturation.
If foundational models saturate regardless of scale, it may be more efficient invest resources into creating direct training data.
4.4 Rich Representations: The Beginnings of a Solution?
(Zhang and Bottou, 2023) propose learning richer representations to capture these missing features. They construct models by concatenating feature representations of individual models and demonstrate significant improvements in accuracy of transfer.
To construct rich representations, they concatenate four copies of ResNet50 trained separately on ImageNet1K, with different random seeds, into an ensemble model (ResNetCat4). They evaluate against a baseline model (ResNet50W2) of the same increased size, trained once on ImageNet1K.
Figure 4: Richer Representations via Concatenation An ensemble model (ResNet50Cat4) is created by concatenating four ResNet models trained separately on ImageNet using different random seed. It is evaluated against a baseline model (ResNet50W2), which is a single ResNet model of the same size, trained once on ImageNet.
Table 2: Supervised transfer learning from ImageNet to iNat18, Cifar100, and Cifar10 using linear probing. The cat4 ensemble concatenates representations of $4$ separately trained networks. The baseline uses a ResNet model of similar size is trained once. These models contain approximately four times the parameters of the original ResNet50.
Direct Transfer via LinProb
Method Arch Params ImageNet1k iNat18 Cifar100 Cifar10
original resnet50 23.5m 75.58 37.91 73.23 90.57
baseline resnet50w2 93.9m 77.58 37.34 72.65 90.86
ensemble cat4$\times$×resnet50 94m 78.15 46.55 78.19 93.09
They find the concatenated ensemble to perform comparably to the baseline on ImageNet1K, but far better when transferred to new datasets, ($+9\%$ on INat18) and ($+5.5\%$ on CIFAR100). Additional comparison to the original ResNet50 reveals that only increasing size of a model (such as the baseline) will worsen performance of transfer: (.$-.5\%$ on INat18), and (.$-.6\%$ on CIFAR100).
The performance gain of the concatenated ensemble can only be attributed to the features recovered by a richer, combined representation. This means that the four concatenated networks learn different subsets of features when trained on identical data shuffled differently.
This also explains why performance on ImageNet1K remains relatively unchanged. Observe that these subsets of features work similarly on ImageNet, that is, each set of features provides the same information to reduce training error. Combining these features marginally improves performance on ImageNet1K ($+2\%$). Hence, they are not learned during training.
Using a similar concatenation approach, (Zhang et al., 2022) demonstrate increased returns in accuracy for Self-Supervised Learning as scale increases from 100M to 1B parameters, as well as on Vision Transformers as scale increases from 100M to 400M parameters (Figure 4). In other words, these rich representations recover crucial features that otherwise would be lost and significantly increases performance accuracy without incurring additional training costs.
Figure 5: Increased Returns in Accuracy as Scale Increases (Left) SSL, 100M to 1B params: The concatenated methods (in red and purple) outperform the baseline (dotted blue curve) on both SWAV trained on unlabeled ImageNet1K (top), and SEER on INSTAGRAM1B (bottom). (Right) ViT, 100M to 400M params: Concatenated representations (purple) outperform the baseline (red) consistently during transfer in both original (top) and modified (bottom) vision transformers (ViT)
4.5 Connections to Multi-Calibration
Missing features fundamentally limits the ability of a model to generalize to new distributions outside the training distribution (OOD). (Wald et al., 2021) finds a link between OOD performance and multi-calibration, introduced by (Hebert-Johnson et al., 2018). Specifically, they argue that calibration leads to better OOD generalization, by creating invariant representations (Arjovsky et al., 2019). (Guy et al., 2025) successfully demonstrate this empirically for in-distribution generalization (IID). However, any remediation requires recovering these missing features.
5 Experiments
5.1 Rich Representations That Cost Nothing
We propose an analogous method to construct the rich representations, inspired by (Zhang et al., 2022), which concatenates networks pretrained for fewer steps. Unlike previous cat experiments that preserve model size (i.e. - space), our method preserves pretraining compute (time). Our findings reveal that the (time-cat) method improves generalization to new datasets while maintaining the cost of pretraining.
5.1.1 Method
Our approach concatenates feature representations from multiple ResNet50 pretrained on ImageNet1k for a shorter duration (Deng et al., 2009), (He et al., 2016).  [^4] Implementation can be found here: https://github.com/facebookresearch/richreps-timecat Each ResNet50 networks is trained identically except using different random seeds. We transfer these concatenated networks using linear probing on new datasets: iNaturalist18, CIFAR-100, CIFAR-10 (Van Horn et al., 2018), (Krizhevsky, 2009).
Our baseline uses a single ResNet50 model trained on ImageNet1k for 4̃50k steps, (reaching 0.65 accuracy on ImageNetv2). We compare how this baseline performs when transferred against several experiments which distribute total pretraining across a number of networks: two ResNet50 networks each trained for 200k steps, four ResNet50 networks each trained for 100k steps, and five ResNet50 networks each trained for 80k steps.
[…]
All ResNet50 models are pretrained using the MLCommons AlgoPerf training framework and the AdamWScheduleFree optimizer (Kasimbeg et al., 2025), (Defazio et al., 2024). The models are transferred using SGD.
5.1.2 Results
Our method yield considerable improvements in performance despite maintaining the same total pretraining computation. By combining more networks that have been pretrained for fewer steps, our experiments produce models which generalize significantly better to new datasets: ($+12.5\%$) on iNaturalist18 and ($+6\%$) on CIFAR-100.
As expected, adding more sets of features does not significantly improve performance ImageNet1K. Since ResNet50 models learn features that provide similar information to reduce training loss, combining them on the training data does not does not yield further improvements.
If concatenating more models trained on fewer steps can yield a significantly better performing model on new datasets, there may be considerable value left on the table.
Table 3: Transfer Performance of Concatenated Models, Each Trained For Fewer Iterations A baseline ResNet50 model (cat1) is trained for 450k steps on ImageNet1k. Each experimental method (catn) creates a models by concatenating multiple (n) ResNet50 networks, each trained independently for 400k/n steps. Overall computational effort of pretraining is maintained.
Pretraining Steps Transfer via LinProb
Method Per ResNet50 Total ImageNet1k Inat18 Cifar100 Cifar10
(BL) cat1 450k 450k 78.4% 33.7% 71.0% 90.4%
cat2 200k 400k 78.5% 40.4% 74.2% 91.2%
cat4 100k 400k 78.3% 42.9% $76.8\%$ 92.7%
cat5 80k 400k 77.8% 46.2% 77.1% 92.8%
6 Limitations
This paper identifies a behavior in deep neural networks, but does not claim to fully discover, solve, or explain this. First, this paper acknowledges prior work on related issues, such as spurious features from imbalanced datasets or poor generalization from foundation models. This paper builds on previous research by using their experimental results as a foundation for its own findings. While it presents rich representations as a potential solution, this paper also recognizes that this solution is incomplete. Instead, it serves as the initial steps towards understanding the problem. Finally, this paper focuses on specific toy examples and empirical evidence to provide insight into when and how this behavior might occur. These results are not comprehensive but provide an early phase of understanding, beginning with the classical case of fine-tuning.
7 Conclusion
This paper explores whether training models across a mixture of tasks provides the necessary features for each component. Our results indicate that supervised transfer methods, however useful, often permanently lose essential features without justification, potentially limiting their ability to generalize effectively. External factors like dataset imbalance or random seed limit what subpopulations a model can learn before training starts. Foundation models cannot alleviate this phenomenon with scale alone. We identify the beginnings of a solution to retrieve these lost features, revealing a fresh problem space with significant potential.
Appendix A Additional Empirical Evidence
This paper introduces a phenomenon called the information saturation bottleneck where models saturate information from encoded features and, consequently, miss key features. This can cause models to perform inconsistently across different subgroups of data.
This bottleneck appears not just in the cases discussed in this paper, but is widespread across the existing machine learning literature. Below are a few additional examples where this issue has been observed.
Biological Foundation Models
Beyond the work in this paper, a number of other works use supervised model to match the performance of larger-scale models.
  • Convolutional method can match the performance of Transformer-based foundation models [Yang et al., 2024]
  • Domain-specific generative model can outperform pretraining for single-cell biology tasks [Kedzierska et al., 2023]
Fairness
  • Fairness treatment of individuals within classification tasks [Dwork et al., 2012]
  • Identifying biased treatment (discrimination on sensitive attributes) in supervised learning [Hardt et al., 2016]
  • Incompatibility between fairness conditions and accuracy of risk scores [Kleinberg et al., 2016]
Robustness
  • Data augmentation for the ImageNet task [Balestriero et al., 2022], [Kirichenko et al., 2023a]
  • Group-DRO on water-bird, CelebA, MultiNLI tasks [Sagawa et al., 2020]

### v3

@@
 Abstract
-Transfer learning is a cornerstone of modern machine learning, promising a way to adapt models pretrained on a broad mix of data to new tasks with minimal new data.
-However, a significant challenge remains in ensuring that transferred features are sufficient to handle unseen datasets, amplified by the difficulty of quantifying whether two tasks are "related".
-To address these challenges, we evaluate model transfer from a pretraining mixture to each of its component tasks, assessing whether pretrained features can match the performance of task-specific direct training.
-We identify a fundamental limitation in deep learning models – an "information saturation bottleneck" – where networks fail to learn new features once they encode similar competing features during training.
-When restricted to learning only a subset of key features during pretraining, models will permanently lose critical features for transfer and perform inconsistently on data distributions, even components of the training mixture.
-Empirical evidence from published studies suggests that this phenomenon is pervasive in deep learning architectures – factors such as data distribution or ordering affect the features that current representation learning methods can learn over time.
-This study suggests that relying solely on large-scale networks may not be as effective as focusing on task-specific training, when available.
-We propose richer feature representations as a potential solution to better generalize across new datasets and, specifically, present existing methods alongside a novel approach, the initial steps towards addressing this challenge.
+Transfer learning is widely used to adapt large pretrained models to new tasks with only a small amount of new data.
+However, a challenge persists – the features from the original task often do not fully cover what is needed for unseen data, especially when the relatedness of tasks is not clear.
+Since deep learning models tend to learn very sparse representations, they retain only the minimal features required for the initial training while discarding potentially ones for downstream transfer.
+A theoretical framework developed in this work demonstrates that such pretraining captures inconsistent aspects of the data distribution, therefore, inducing transfer bias.
+To address this limitation, we propose an inexpensive ensembling strategy that aggregates multiple models to generate richer feature representations.
+On ResNet, this approach yields a $9\%$ improvement in transfer accuracy without incurring extra pretraining cost.
+We also present empirical evidence from a range of deep learning studies, confirming that the phenomenon is pervasive across modern deep learning architectures.
+These results suggests that relying solely on large pretrained networks is not always the most effective way to improve model generalization.
+Instead, fostering richer, more diverse representations – e.g.
+- through model ensembles – can substantially enhance transfer learning performance.
 1 Introduction
 Transfer learning has become a cornerstone of modern machine learning and artificial intelligence (Bottou, 2011; Collobert et al., 2011; Oquab et al., 2014; Bommasani et al., 2021).
-In its simplest form – supervised transfer learning, a relatively large deep learning network is pre-trained on a large and diverse dataset composed of cheap examples that are somehow related to the task of interest.
-In a second phase, the network is adapted using a much smaller dataset composed of examples that directly illustrate the task of interest.
-This second dataset is often too small to train a network of that size from scratch (Li et al., 2017).
-It is hoped that the representations constructed during the pretraining phase will be useful during the adaptation phase.
-It remains an significant challenge to ensure that transferred features are sufficient to handle new, unseen datasets.
-A pretrained network that is missing crucial feature information may not perform on par with direct learning, that is, a network with all the features we could obtain by training directly on enough data to learn a target task.
-Foundation model folklore suggests that, by pretraining these enormous models on "everything", they would learn "all the features (abilities) we could ever want." Should we invest into building ever larger all-purpose foundational models, or into collecting tasks-specific datasets for training smaller specialized models?
-Positive findings can only validate transfer from a particular foundational model to a particular target task.
-We cannot say more because we do not know how to even state that the target task is related to the pre-trained model “abilities”.
-Instead, we provide a negative result by exposing shortcomings that appear under very favorable conditions.
-We assume that the pre-training data is drawn from a mixture of distribution that includes the target task as a component, and we show that we cannot expect that an adapted model will always match a directly trained model.
-If we cannot always match direct training when adapting a model pre-trained on a mix that includes the target task, we cannot expect it either when we know even less about the pre-training data.
-We uncover an information saturation bottleneck that occurs when models cannot learn new features after they have encoded similar competing features.
-When networks are restricted to are restricted to learning only a subset of key features during pretraining, they lose crucial feature information that is essential to effectively adapt to new tasks, or even ones they have seen before.
-These lost features are nearly impossible to recover by simply fine-tuning on limited data.
-Although this phenomenon was not previously recognized as such, we argue that it is pervasive in deep learning models.
-We analyze this phenomenon rigorously through a simple theoretical counterexample, and discuss several published papers whose empirical findings can be explained by a consistent loss of features that would be useful for transfer.
-Additionally, factors such as the distribution imbalances or shuffling order of the training can determine which features a model can learn over time.
-Finally, we propose richer feature representations as a potential solution to improve generalization across new datasets and present existing methods alongside novel approaches which represent the beginning of addressing this challenge.
-Section 2 presents a formal framework to assess effectiveness of transfer.
-Section 3 demonstrates through a simple example, that even in optimal conditions, a deep network might miss key features.
-Section 4 reveals a fundamental bottleneck of deep networks, supported by empirical evidence across existing studies.
-It also introduces "rich representations" as a potential solution.
-Section 5 presents a method of constructing rich representations that improves performance when transferred to unseen distributions, at no additional pretraining cost.
-Why study Mixture Distributions?
-This paper examines how real-world pretraining works by studying complicated datasets with many imbalanced subgroups.
-We use a mathematical model called mixture distribution to represent these complex datasets as a single cohesive structure made up of multiple probability distributions.
-Mixture of Experts
-This is distinct from the neural network architecture known as "mixture of experts" (MoEs), where specialized "expert" models learn different parts of a task.
-When training on real-world datasets with many components, each "expert" may focus on smaller number tasks, but still be bound by the same limitations.
+In its simplest form–supervised transfer learning–a large neural network is first trained on inexpensive and diverse dataset that is related in some way to the target task.
+The network is then fine-tuned on a much smaller, task-specific dataset.
+Since the fine-tuning dataset is traditionally too small too small to train a network of that size from scratch, the hope is that the representations learned during the pretraining phase will be useful during the adaptation phase (Li et al., 2017).
+A key challenge is ensure that transferred features are sufficient for unseen data.
+When pre-training data is unrelated to the target task, the adapted model can underperform compared with a model trained directly on ample task-specific data.
+The folklore around foundation model suggests that training an enormous model on "everything" will give it essentially all useful features Bommasani et al.
+(2021).
+We ask a more modest question: if a model is pretrained on a mixture of data that includes the target task, can the fine-tuned model match the performance of a model trained directly on a large, task-specific dataset?
+Answering this question informs whether researchers should invest in ever larger, all-purpose foundation models or in collecting task-specific datasets for training smaller specialized models.
+Our findings are negative : even under favorable conditions–a pretraining mixture that contains the target task–a fine-tuned model will still fall short of a model trained directly on ample task data.
+In other words, pretraining on mixed data does not guarantee that fine-tuning will achieve parity with task-specific training.
+If a model pretrained on mixed data that includes the target task cannot consistently match a model trained directly on that task, then we should not expect good performance when we have little knowledge of how the pretraining data relates to the target task.
+We uncover a fundamental bottleneck caused by the sparsity bias of deep learning networks.
+During pretraining, a model learns only the minimal set of features required for that task, discarding other features that may be essential for downstream problems.
+As a result, crucial information is lost and cannot be recovered by simply fine-tuning on a small target dataset – even when the target task is represented in the pretraining data.
+Once a feature is encoded, a network often fails to learn competing features, which hampers adaptation to new or previously seen tasks.
+Although previously overlooked, this phenomenon is pervasive.
+We illustrate it with a simple theoretical counterexample and discuss empirical studies that reveal consistent feature loss.
+Factors such as class-imbalance and data-shuffling order further influence which features are retained.
+To address this limitation, we propose an inexpensive strategy that improves generalization to unseen distributions without additional cost.
+For example, on ResNet50, this approach yields a $9\%$ improvement in transfer accuracy without incurring extra pretraining cost.
+The paper is organized as follows:
+  • Section 2 formalizes the effectiveness of transfer.
+  • Section 3 presents the counterexample.
+  • Section 4 discusses the bottleneck, supports it with empirical evidence, and introduces rich representations as a solution.
+  • Section 5 describes a method for constructing such representations and demonstrates its benefits on unseen data..
 2 Problem Statement
-Consider a set of distributions $\{P^{[j]}(X,Y)\}_{j}$ representing data from different populations, and a mixture distribution $P^{[\mathrm{mix}]}(X,Y)$ formed by a weighted sum of these subdistributions.
-[^1] Note: This mixture is non-trivial, meaning that all component distributions are present.
-$P^{[\mathrm{mix}]}(X,Y)=\sum_{j}\lambda_{j}\cdot P^{[j]}(X,Y)\quad\text{where}~{}\lambda_{j}>0,~{}\sum_{j}\lambda_{j}=1.$
-Suppose that we pretrain a network on mixture $P^{[\mathrm{mix}]}$.
-Let $P^{[i]}$ denote a fixed arbitrary subdistribution.
-Assuming that we have enough data to directly train a network on $P^{[i]}$, we ask
-Question: Do features pretrained using $P^{[\mathrm{mix}]}$ work nearly as well as features learned directly for target $P^{[i]}?$
-2.1 Two Ways to Train
-Consider a neural network $f$ of the following form
-$f(X;\theta,\bm{\gamma})=\sum_{j}\,\underbrace{\varphi_{j}(X;\theta)}_{\text{feature extractor}}\cdot\underbrace{\gamma_{j}}_{\text{classifier}}$ (1)
-This network has two main components.
-The feature extraction layers ($\varphi_{j}(X;\theta)$) initially transform the raw input $X$ into real-valued features using parameters $\theta.$ Furthermore, these transformations may be non-linear, enabling the extraction of complex signals from the input data.
-The linear classification layer ($\gamma_{j}$) combines linearly the extracted features to produce the final prediction.
-[^2] We explain in the following section why it is sufficient to study the linear case.
+Let $\{P^{[j]}\}_{j}$ denote a set of probability distributions over (X,Y), each corresponding to data from a distinct population. Consider a non-trivial mixture
+$P^{[\mathrm{mix}]}(X,Y)=\sum_{j=1}^{n}\lambda_{j}P^{[j]}(X,Y),\quad\lambda_{j}>0,~\sum_{j=1}^{n}\lambda_{j}=1.$
+Assume a neural network is pretrained on samples drawn from $P^{[\mathrm{mix}]}$.
+Let $P^{[i]}$ be an arbitrary target distribution from the same family.
+If we have enough data, we also could train a network directly on $P^{[i]}$.
+Question: Do the feature representations learned from pretraining on $P^{[\mathrm{mix}]}$ achieve performance on $P^{[i]}$ that is comparable to a networks directly trained on $P^{[i]}?$
+In other words, when the mixture contains the target distribution, can transfer learning from the mixture match the efficacy of task-specific training
+2.1 Two Ways to Train a Neural Network
+We study a network of the form
+$f(X;\theta,\bm{\gamma})=\gamma^{T}\ \varphi(X;\theta)$ (1)
+Where $\varphi(\cdot;\theta)$ is a feature extractor (e.g. - a stack of convolutional neural networks or transformer layers) and $\gamma$ is a linear classifier.
+The extractor transforms the raw input $X$ into real-valued features using parameters $\theta$; the classifier produces the final prediction.
 There are two ways to train such a model for distribution $P^{[i]}$ (Figure 1)
-  1. 1.
-    The model could be directly trained on the target distribution $P^{[i]}.$ In this case, the model learns both feature parameters $\theta$ and classifier weights $\bm{\gamma}$ optimized for the target distribution $P^{[i]}$.
-This yields the final trained model $f(\cdot\ ;\theta^{[i]},\bm{\gamma}^{[i]})$.
-  2. 2.
-    Alternatively, we could train a model through transfer learning (via linear probing).Footnote 2 The model is pretrained on a mixture distribution $P^{[\mathrm{mix}]}$.
-The feature parameters $\theta^{[\mathrm{mix}]}$ are frozen and only the linear classifier weights $\bm{\gamma}$ are fine-tuned on $P^{[i]}$.
-This is called linear probing.
-The two approaches primarily differ in how they learn features.
-In direct training, a model is capable of capturing any representable feature from the target distribution $P^{[i]}$.
-However, in transfer learning, the model is restricted to selecting features based on a different pretraining task $P^{[i]}$, which includes $P^{[i]}$ as a component among many.
-Under the constraints of transfer learning, it is unclear whether models pretrained on broad mixtures retain sufficient features to match the performance task-specific training.
-To assess their effectiveness in preserving performance, we explore whether the direct training solutions $f(X;\theta^{[i]},\gamma^{[i]})$ of any component $P^{[i]}$ can be matched (or approximated) by linearly combining features $\varphi_{j}(X;\theta^{[\mathrm{mix}]})$ from pretraining on the mixture.
-We assume access to unlimited training data, which is important for effective direct training.
-Theoretically, we ask whether the function space spanned by features $\varphi_{j}(X;\theta^{[\mathrm{mix}]})$ contains the direct training solutions $f(\dots,\theta^{[i]},\gamma^{[i]})$ trained on unlimited data.
-Figure 1: Two ways to train a classifier for $P^{[i]}$: directly from $P^{[i]}$ or transferred from $P^{[\mathrm{mix}]}$.
-2.1.1 Studying the Linear Case is Sufficient
-Linear Probing:
-We focus on the scenario of transfer learning via linear probing, where only the final classification layer is modified during fine-tuning. Thus, our results directly apply to linear probing.
-General Fine-Tuning:
-Our findings extend beyond linear probing to more common case where all network layers are adjusted. This can be written as a particular instance within our framework.
+Direct training on a target distribution
+Given a data distribution $P^{[i]}$, we could learn both the extractor and classifier from scratch
+$(\theta^{[i]},\gamma^{[i]})=\arg\min_{\theta,\gamma}\mathbb{E}_{(X,Y\sim P^{[i]})}[\ell(f(X;\theta,\gamma),Y)]$
+The resulting model $f(\cdot;\theta^{[i]},\gamma^{[i]})$ is fully adapted to $P.$
+Transfer learning via linear probing
+We first pretrain the extractor on a mixture distribution
+$P^{[\mathrm{mix}]}(X,Y)=\sum_{j}\lambda_{j}P_{j}(X,Y),\quad\lambda_{j}>0,\sum_{j}\lambda_{j}=1$
+obtaining parameters $\theta^{[\mathrm{mix}]}$.
+These parameters are frozen and only the classifier is trained on $P^{[i]}$:
+$\gamma^{[i]}_{\quad lin}=\arg\min_{\gamma}\mathbb{E}_{(X,Y)\sim P^{[i]}}[\ell(f(X;\theta^{[\mathrm{mix}]},\gamma),Y)].$
+[…]
+The Central Question
+With unlimited data for direct training, can the performance of the directly trained model $f(\cdot;\theta^{[i]},\gamma^{[i]})$ be matched or approximated by the model that uses the pretrained features $\varphi(\cdot;\theta^{[\mathrm{mix}]})$ and only trains a linear classifier?
+In other words
+$\mathbb{E}_{(X,Y)\sim P^{[i]}}[\ell(f(X;\theta^{[i]},\gamma^{[i]}),Y)]$
+$\quad\to(?)\mathbb{E}_{(X,Y\sim P^{[i]})}[\ell(f(X;\theta^{[\mathrm{mix}]},\gamma^{[i]}_{\quad lin}),Y)]$
+Equivalently, does the function space spanned by the pretrained mixture $\{\gamma^{T}\varphi(\cdot;\theta^{[\mathrm{mix}]})\}$ contain the solutions $\{f(\cdot;\theta_{i},\gamma_{i})\}$ obtained by training directly on the target distribution $P^{[i]}$?
+Answering these questions tells us whether pretraining on a broad mixture that includes the target task suffices to capture all features needed for optimal performance on that task.
+2.2 Data Assumptions
+We take the following assumptions for a realistic model example.
+  • Unlimited pre‑training data – We assume an infinite stream of samples from the pre‑training distribution. Even with unlimited data, a model can still miss essential features.
+  • Finite fine-tuning data – Transfer learning is applied to a small target set, which limits the adaptation of learned features.
+  • Infinite data for linear probing – Fine tuned weights stay close to the pre-training weights (see (Radiya-Dixit and Wang, 2020)), so fine-tuning can be treated as linear probing on the NTK features.
+We assume unlimited training data, which ensures that a model directly trained on $P^{[i]}$ can achieve perfect generalization.
+Figure 1: Classifier trained directly from $P^{[i]}$ versus transferred from the mixture $P^{[\mathrm{mix}]}$.
+2.3 Linear probing is sufficient
+We first focus on linear probing, in which only the final classification layer is fine-tuned.
+Our analysis directly applies to this common scenario and it also covers fine-tuning because the weight updates can be expressed as a linear combination in the pre-trained features space.
+Fine-Tuning as a linear combination
+Let $f(X;\mathbf{w}_{{}^{[\mathrm{mix}]}})$ be the pre-trained network.
+After fine-tuning on $P^{[i]}$, the new weights $\mathbf{w}^{[i]}$ remain close to the $w^{[\mathrm{mix}]}$.
+A first order Taylor expansion gives
 Consider the general case of fine-tuning a pre-trained neural network $f(X;\mathbf{w}^{[\mathrm{mix}]})$ that updates all of its weights for target $P^{[i]}$.
@@
 […]
-Observe that $f(X;\mathbf{w}^{[i]})$ is a linear combination of ${\partial f}/{\partial w_{j}}(X;\mathbf{w}^{[\mathrm{mix}]})$, which are exactly the neural tangent kernel (NTK) features (Jacot et al., 2018).
-The fine-tuning operation effectively modifies only the linear coefficients $(w_{j}^{[i]}-w_{j}^{[\mathrm{mix}]})$.
-Thus, this is an instance of our abstract problem statement, in the space of these NTK features.
-Non-Linear Classifier:
-It is also common to fine-tune multiple non-linear classification layers of a network. The previous argument also applies when we modify some, not all, weights
-2.1.2 Networks with Non-Zero Training Error
-When training neural networks, it is common practice to optimize the model until the training error reaches zero. However, we cannot learn useful information from studying these.
-Transfer has historically been used when there is insufficient data to directly learn a task.
-A network may achieve zero training error without generalizing perfectly to the entire distribution.
-When not all features have been learned, transfer to a subdistribution indicates how a network may generalize.
-Our problem setup makes the assumption of unlimited training data.
-When a model has trained on everything, zero training error translates to zero test error.
-When a network already performs perfectly on all subdistributions, transfer to a component tells us nothing about how it may generalize.
-2.1.3 Covariance
+$\quad\underbrace{\frac{\partial f}{\partial w_{j}}(X;\mathbf{w}^{[\mathrm{mix}]})}_{\varphi_{j}(X;\ \theta^{[\mathrm{mix}]})}~\times~\sum_{j}\underbrace{(w_{j}^{[i]}-w^{[\mathrm{mix}]}_{j})}_{\gamma_{j}}.$
+[…]
+2.4 Models with non‑zero training error
+Training a network until zero training error on the entire mixture would make it learn every feature, so transfer would be trivial.
+Such a case does not often appear in real-world settings.
+To study the limits of transfer in a realistic setting, we therefore consider networks that still exhibit non-zero training error after pretraining.
+In realistic settings – especially with mixture distributions – a network may overfit on high-weight components and under-fit on low-weight ones, leaving some components with non-zero test error.
+2.5 Covariance
+A feature that is correlated with the label in any component of the mixture remains correlated under almost all mixtures.
 Proposition 2.1.
-Suppose we find a feature $\varphi(X)$ correlated output $Y$ under the subdistribution $P^{[i]}$ Then, $\varphi(X)$ is correlated with $Y$ under any mixture $P^{[\mathrm{mix}]}=\sum_{j}\lambda_{j}P^{[j]}$, except maybe for a measure zero set of mixture coefficients $(\lambda_{j})$.
-$\begin{split}\forall\neq 0\quad\implies\mathrm{Cor}_{P^{[\mathrm{mix}]};\lambda}(\varphi(X),Y)\neq 0\quad\text{for }\text{ a.e. }\mathbf{\lambda}\end{split}$
+If $\mathrm{Cor}_{P^{[i]}}[\varphi(X),Y]\neq 0$, for some component $P^{[i]}$, then
+$\mathrm{Cor}_{P^{[\mathrm{mix}]}}[\varphi(X),Y]\neq 0$
+for almost all mixtures $P^{[\mathrm{mix}]}$; only a measure-zero set of mixture proportions can cancel out the correlation.
 Proof.
-Assume that this is not the case. Two variables are correlated if and only if they have nonzero covariance. Therefore, mixture coefficients $\lambda_{i}$ must satisfy the equation:
-$0=\mathrm{Cov}^{[\mathrm{mix}]}(Y,\varphi(X))=\sum_{j}\lambda_{j}\mathrm{Cov}^{[j]}(Y,\varphi(X)).$
-Under our assumption, at least one of the covariances $\text{Cov}^{[j]}$ is nonzero.
-Thus, this equation represents a measure zero subset of mixture coefficients.
-Outside of this subset, $\mathrm{Cov}^{[\mathrm{mix}]}(Y,\varphi(X))\neq 0$.
-∎
-Intuitively, the remaining components must exactly cancel out the positive covariance of $P^{[i]}$, which only occurs for precisely-balanced mixture coefficients.
-This suggests that a feature interesting to a target $P^{[i]}$ (i.e.
-- correlated with Y) is also interesting to the pretraining distribution $P^{[\mathrm{mix}]}$.
-However, a nonzero covariance alone does not mean that this feature will be part of the solution $f(\theta^{[\mathrm{mix}]},\gamma^{[\mathrm{mix}]})$ of the pre-training problem (useful).
-If the information $\varphi(X)$ provides about $Y$ has already been introduced by existing features in the solution, it is no longer useful.
-This is consequential when the training setup is biased towards learning sparse representations.
+Assume the contrary. First, two variables are correlated if and only if they have nonzero covariance. Therefore, mixture coefficients $(\lambda_{j})$ would have to satisfy
+$0=\sum_{j}\lambda_{j}\mathrm{Cov}^{[j]}(Y,\varphi(X)).$
+a linear equation that is satisfied only on a measure-zero set of $\lambda_{i}$. Hence, the proposition holds. ∎
+Thus, interestingness (non-zero correlation) transfers almost always, but correlation alone does not guarantee that a feature is learned in the solution.
 Sparsity in deep networks
-Stochastic gradient learning algorithms in deep learning models often have an implicit sparsity bias (Gunasekar et al., 2017; Andriushchenko et al., 2023), a likely contributor to the effectiveness of deep learning.
-In an oversimplified account of the analysis of Andriushchenko et al., when using stochastic gradient descent with large step sizes, features are added and removed somewhat randomly, but those that contribute to reducing the training error tend to stick around.
-Depending on the random order in which features are discovered by the training algorithm, it may construct different sparse solutions.
+Stochastic gradient learning algorithms in deep learning models often have an implicit sparsity bias (Gunasekar et al., 2017; Andriushchenko et al., 2023), which likely contributes to the effectiveness of deep learning.
+In an oversimplification of Andriushchenko et al., when using SGD uses a large step sizes, it retains only features that reduce the training error; the random order of feature discovery can lead to different sparse models.
 Many practical regularization tricks such as early-stopping or weight decay have a similar effect.
@@
 3 Counterexample
-This study explores whether the features learned from mixture distribution perform as well across component tasks $P^{[i]}$ as features learned specifically for each task.
-We introduce a simple counterexample demonstrating that this desired outcome does not hold.
-Example Setup
-Imagine a feature extractor which can learn only two feature extraction functions $\Phi(X)=(\varphi_{1}(X),\varphi_{2}(X))$.
-Depending on parameters $\theta,$ the space spanned by the selected features can have zero, one, or two dimensions.
-Consider four individual distributions in this space, illustrated in Figure 3:
+We show that training on a mixture does not guarantee that learned features are as useful for each component task.
+Consider a feature extractor that can produce only two real-valued functions
+$\Phi(X)=(\varphi_{1}(X),\varphi_{2}(X))$
+. Depending on parameters $\theta,$ the extractor may use one, two, or none of these features.
+Four toy distributions are defined on the two-dimensional space of $\Phi(x)$, illustrated in Figure 3:
   • $P^{[1]}[\>\Phi(X){=}(+1,0),\>Y{=}+1\>]=\tfrac{1}{2}$ and $P^{[1]}[\>\Phi(X){=}(0,\pm 1),\>Y{=}-1\>]=\tfrac{1}{4}$.
@@
   • $P^{[4]}[\>\Phi(X){=}(0,-1),\>Y{=}-1\>]=\tfrac{1}{2}$ and $P^{[4]}[\>\Phi(X){=}(\pm 1,0),\>Y{=}+1\>]=\tfrac{1}{4}$.
-Each distribution is defined across three distinct points: two points with probability $1/4$, and the third point with probability $1/2$.
-Combining these forms a mixture $P^{[\mathrm{mix}]},$ supported over four points (Figure 3).
-$P^{[\mathrm{mix}]}=\lambda_{1}\cdot P^{[1]}+\lambda_{2}\cdot P^{[2]}+\lambda_{3}\cdot P^{[3]}+\lambda_{4}\cdot P^{[4]}\quad\text{where}~{}\lambda_{i}>0,~{}\sum_{i=1}\lambda_{i}=1$
-A distribution $P^{[\mathrm{mix}]}$ with non-zero mixture coefficients $\lambda_{i}$, therefore, offers positive examples at $\Phi(X)=(\pm 1,0)$ and negative examples at $\Phi(X)=(0,\pm 1)$.
+Each distribution places a point of mass $1/2$ on a positive example and two points of mass $1/4$ on negative examples. The mixture
+$P^{[\mathrm{mix}]}=\sum_{i=1}^{4}\lambda_{i}P_{i},\quad\lambda_{i}>0,~\sum_{i=1}\lambda_{i}=1$
+has support on four points: the two positive points $\pm 1,0$} and the two negative points $(0,\pm 1)$. See Figure 3.
 The Optimal Classifier
-Suppose that we optimize binary classifiers for each of the four distributions.
-The training error of each $P^{[i]}$ is minimized using one specific feature: $\varphi_{1}(X)$ for $P^{[1]}$ and $P^{[2]}$, $\varphi_{1}(X)$ for $P^{[3]}$ and $P^{[4]}$.
-Now, consider a binary classifier optimized for $P^{[\mathrm{mix}]}.$ Since its four points cannot be linearly separated, the training error is minimized by misclassifying the least-weighted point and correctly classifying the other three.
-Depending on which point is ignored, the optimal classifier is one of the four solutions in Figure 3
+For each $P_{i}$, the optimal binary classifier uses a single feature $\varphi_{1}$ for $P^{[1]}$ and $P^{[2]}$ and $\varphi_{2}$ for $P^{[3]}$ and $P^{[4]}$.
+The mixture cannot be linearly separated; the optimal classifier misclassifies the lowest-weighted point and correctly classifies the other three.
+Depending on which point is ignored, the solution uses either $\varphi_{1}$ or $\varphi_{2}$.
+Since deep learning networks with sparsity bias learn only one useful feature, the mixture-trained model will choose either $\varphi_{1}$ or $\varphi_{2}$.
+Consequently, it can correctly classify only two of the four sub-tasks; if it learns $\varphi_{1}$, it fails on $P3$ and $P_{4}$.
+If it learns $\varphi_{2}$, it fails on $P_{1}$ and $P_{2}$, as in Figure 3/
+Thus, even in this favorable setting, pretraining on a mixture can miss essential features for its subtasks.
 Figure 2: Four subdistributions and a combined mixture distribution are represented by red points (labeled $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+1$) and blue points (labeled $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1$) Each subdistribution includes three points; two points contain an equal number of examples, while the third point contains twice as many.
@@
 Each classifier uses a dotted line to separate red points (labeled $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+1$) from the blue points (labeled $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1$).
-Sincef points are not linearly separable, an optimal classifier ignores the least-weighted point (smallest) and classifies the remaining three points.
+Since the points are not linearly separable, an optimal classifier ignores the least-weighted point (smallest) and classifies the remaining three points.
 Interestingly, this optimal classifier can be sparsely represented using just one feature, either $\varphi_{1}(X)$ or $\varphi_{2}(X),$ but cannot classify two of the four subdistributions.
-Observe that the optimal classifier $P^{[\mathrm{mix}]}$ can be sparsely represented using just one of the two features, determined by the balance of the mixture.
-Hence, a deep neural network which exhibits sparsity bias will represent this optimal solution by learning just one feature.
-In other words, the network learns a feature space that does not contain the optimal solutions for two of its four subdistributions.
-Specifically, when it learns only $\varphi_{1}$, it cannot correctly classify $P^{[3]}$ and $P^{[4]}$; when it learns only $\varphi_{2}$, it cannot correctly classify $P^{[1]}$ and $P^{[2]}$.
-[^3] Some mixtures have multiple clusters of equal minimal weight and, therefore, multiple optimal solutions.
-However, this is a set of negligible probability.
+ [^1] Some mixtures have multiple clusters of equal minimal weight and, therefore, multiple optimal solutions. However, this is a set of negligible probability.
 We find that training a deep network (with sparsity bias) on the mixture distribution yields a sparse representation which can classify only half of its subdistributions.
 Even in this advantageous setup, a model pretrained on a mixture may miss important features for its subtasks.
-3.1 Assumptions
-In this counterexample, we study the training error as an indicator of how well a model can adapt to its subdistributions. We make the following assumptions
-Assumption 1: A Model Will Learn These Features(Papyan et al., 2020) finds that a neural network learns representations which "collapse" into the problem space features, $\varphi_{1}(X)$, $\varphi_{2}(X)$
-Assumption 2: We Cannot Study Non-Linear Decision Boundaries A network with a non-linear decision boundary could learn both features and optimally classify all subdistributions.
-However, we are not interested in studying solutions with zero training error (see Section 2.1.2).
-4 True in Practice?
-Our simple counterexample illustrates how a model may struggle to learn crucial features during training, which creates biased performance on subtasks of the training distribution.
-When an optimally-trained model can miss a crucial feature in such a simple setup, it suggests that deep networks may frequently miss important feature when trained on more complex datasets.
-In fact, the results of many already-published papers could be explained by this phenomenon.
-Instead of running redundant experiments, we describe this limitation and explore existing experiments below.
-4.1 An Information Saturation Bottleneck
-We identify a fundamental limitation of deep learning models that arises from their sparsity bias – an information saturation bottleneck.
-These networks struggle to learn new features when they have encoded similar competing features during training.
-When networks are restricted to learning a subset of key features during pretraining, they permanently lose critical features needed for effective transfer, leading to inconsistent performance on data distributions–even components of the pretraining mixture.
-The previous counterexample demonstrates that balance of a mixture may determine which features a deep network will learn over time.
-We provide empirical evidence for this and identify additional factors which influence the subset of features learned by a network.
-4.2 Spurious Features
+3.1 Assumptions of the Counterexample
+In this counterexample, use the training error as a proxy for how well a model can later adapt to each subdistributions. The analysis relies on two simplifying conditions
+Assumption 1: Feature learning collapses to the problem space.(Papyan et al., 2020) finds that a neural network learn representations which "collapse" into the problem space features; in our toy example, these are the two basis functions $\varphi_{1}(X)$, $\varphi_{2}(X)$ .
+Assumption 2: We Restrict Attention to Non-Linear Decision Boundaries This type of data (XOR) may realistically appear in the NTK domain of a learned model.
+When fine-tuning the last layer features of this model on a smaller target set, this mirrors linear probing of the NTK features (as explained L108-117).
+There is no way to fine tune or introduce non-linearities within the NTK feature space to overcome this limitation.
+4 Evidence from Existing Papers
+Our counterexample shows that a model can fail to learn essential features, leading to biased performance on sub‑tasks of the training distribution.
+Even when optimally trained, a network may omit a critical feature in this toy setting—suggesting that deep networks often overlook important features when applied to more complex data.
+4.1 Hypothesis: Sparsity Bias Limits Feature Learning
+We argue that deep neural networks possess an inherent sparsity bias–arising from SGD, weight decay, and early stopping–that restricts the set of features that they can learn.
+When a network has encoded a feature that is similar to a new one, the optimizer tends to keep the existing feature and discard the new one.
+Consequently, only a subset of the truly informative features is retained, and critical features that would help later transfer are lost.
+In our counterexample this manifests as inconsistent performance on different components of the pretraining distribution.
+The order in which features are learned depends on arbitrary factors such as data shuffling and class‑balance, which introduces a bias in the resulting representation.
+This perspective explains several previously reported phenomena; rather than repeating experiments, we describe the limitation and survey relevant work below.
+4.2 Example - Spurious Features
 Table 1: Empirical Evidence (Left) Performance gap after recovering core features with transfer (Right) The simple supervised models trained by their supervised ’DASHA’ workflow outperform eight well Genomic Foundation Models (GFMs) on the diverse NT benchmark (Xu et al., 2025)
@@
 Model Model Pretraining Average Average Mean Median
-Size Base-Pairs Score $\uparrow$ Rank $\downarrow$ %Imp.$\uparrow$ %Imp.$\uparrow$
+Size Base-Pairs Score $\uparrow$ Rank $\downarrow$ $\uparrow Imp.\uparrow$
 Enformer 252M 4B 0.569 11.86 27.73 27.91
@@
 In this example of the bottleneck, pretraining on an imbalanced mixture produces misleading spurious features, which prevent the network from learning core features which are essential for prediction.
-4.3 Genomic Foundation Models
+4.3 Example - Genomic Foundation Models
 (Xu et al., 2025) find that simple supervised CNNs can be easily trained to match the performance of transformer-based genomic foundation models (GFMs).
@@
 If foundational models saturate regardless of scale, it may be more efficient invest resources into creating direct training data.
-4.4 Rich Representations: The Beginnings of a Solution?
+4.4 Example - Rich Representations: The Beginnings of a Solution?
 (Zhang and Bottou, 2023) propose learning richer representations to capture these missing features.
-They construct models by concatenating feature representations of individual models and demonstrate significant improvements in accuracy of transfer.
+They concatenating feature representations of individual models and demonstrate significant improvements in accuracy of transfer.
 To construct rich representations, they concatenate four copies of ResNet50 trained separately on ImageNet1K, with different random seeds, into an ensemble model (ResNetCat4).
@@
 Table 2: Supervised transfer learning from ImageNet to iNat18, Cifar100, and Cifar10 using linear probing.
-The cat4 ensemble concatenates representations of $4$ separately trained networks.
+The cat4 ensemble concatenates representations of 44 separately trained networks.
 The baseline uses a ResNet model of similar size is trained once.
@@
 baseline resnet50w2 93.9m 77.58 37.34 72.65 90.86
-ensemble cat4$\times$×resnet50 94m 78.15 46.55 78.19 93.09
-They find the concatenated ensemble to perform comparably to the baseline on ImageNet1K, but far better when transferred to new datasets, ($+9\%$ on INat18) and ($+5.5\%$ on CIFAR100).
-Additional comparison to the original ResNet50 reveals that only increasing size of a model (such as the baseline) will worsen performance of transfer: (.$-.5\%$ on INat18), and (.$-.6\%$ on CIFAR100).
+[…]
+They find the concatenated ensemble to perform comparably to the baseline on ImageNet1K, but far better when transferred to new datasets, ($9\%$ on INat18) and ($5.5\%$ on CIFAR100).
+Additional comparison to the original ResNet50 reveals that only increasing size of a model (such as the baseline) will worsen performance of transfer: ($5\%$ on INat18), and ($6\%$ on CIFAR100).
 The performance gain of the concatenated ensemble can only be attributed to the features recovered by a richer, combined representation.
 This means that the four concatenated networks learn different subsets of features when trained on identical data shuffled differently.
-This also explains why performance on ImageNet1K remains relatively unchanged.
+This also explains why simply increasing model size without also increasing model richness produces similar or slightly worse results.
 Observe that these subsets of features work similarly on ImageNet, that is, each set of features provides the same information to reduce training error.
-Combining these features marginally improves performance on ImageNet1K ($+2\%$).
+Combining these features marginally improves performance on ImageNet1K ($2\%$).
 Hence, they are not learned during training.
@@
 (Right) ViT, 100M to 400M params: Concatenated representations (purple) outperform the baseline (red) consistently during transfer in both original (top) and modified (bottom) vision transformers (ViT)
-4.5 Connections to Multi-Calibration
-Missing features fundamentally limits the ability of a model to generalize to new distributions outside the training distribution (OOD).
-(Wald et al., 2021) finds a link between OOD performance and multi-calibration, introduced by (Hebert-Johnson et al., 2018).
-Specifically, they argue that calibration leads to better OOD generalization, by creating invariant representations (Arjovsky et al., 2019).
-(Guy et al., 2025) successfully demonstrate this empirically for in-distribution generalization (IID).
-However, any remediation requires recovering these missing features.
-5 Experiments
-5.1 Rich Representations That Cost Nothing
-We propose an analogous method to construct the rich representations, inspired by (Zhang et al., 2022), which concatenates networks pretrained for fewer steps.
-Unlike previous cat experiments that preserve model size (i.e.
-- space), our method preserves pretraining compute (time).
-Our findings reveal that the (time-cat) method improves generalization to new datasets while maintaining the cost of pretraining.
-5.1.1 Method
-Our approach concatenates feature representations from multiple ResNet50 pretrained on ImageNet1k for a shorter duration (Deng et al., 2009), (He et al., 2016).
-[^4] Implementation can be found here: https://github.com/facebookresearch/richreps-timecat Each ResNet50 networks is trained identically except using different random seeds.
-We transfer these concatenated networks using linear probing on new datasets: iNaturalist18, CIFAR-100, CIFAR-10 (Van Horn et al., 2018), (Krizhevsky, 2009).
-Our baseline uses a single ResNet50 model trained on ImageNet1k for 4̃50k steps, (reaching 0.65 accuracy on ImageNetv2).
-We compare how this baseline performs when transferred against several experiments which distribute total pretraining across a number of networks: two ResNet50 networks each trained for 200k steps, four ResNet50 networks each trained for 100k steps, and five ResNet50 networks each trained for 80k steps.
-[…]
-All ResNet50 models are pretrained using the MLCommons AlgoPerf training framework and the AdamWScheduleFree optimizer (Kasimbeg et al., 2025), (Defazio et al., 2024).
-The models are transferred using SGD.
+5 Our New Experiments
+5.1 Improved Transfer at Fixed Cost
+We propose a method to improve transfer performance with fixed total compute.
+Our method creates rich representations by ensembling, following the method of (Zhang et al., 2022) to concatenate networks pretrained for fewer steps.
+In contrast to (Zhang et al., 2022)’s experiments which preserve model parameters (space), our method fixes total compute (time). We call our method (time-concat).
+5.1.1 Methodology
+Total pretraining time is fixed at 450k iterations per ResNet50 model on 8xH100 GPUs on ImageNet1k (Deng et al., 2009), (He et al., 2016).
+[^2] Implementation can be found here: https://github.com/facebookresearch/richreps-timecat All pretraining uses the AdamWScheduleFree Optimizer and builds on the MLCommons AlgoPerf Training Framework (Kasimbeg et al., 2025), (Defazio et al., 2024).
+  • Baseline Model: A single ResNet50 model fully-trained on ImageNet1k.
+  • Ensemble Model: Multiple ResNet50 models trained independently for shorter periods (different seeds, then ensembled).
+For example, an ensemble could contain four models trained for $25\%$ of the total time or two models trained for $50\%$ of the total time.
+Each ResNet50 network is trained identically except using different random seeds. We transfer these concatenated networks using linear probing on new datasets.
+5.1.2 Evaluation
+Models are transferred to datasets via linear probing (fine-tuning only the final classification layer) and evaluated on new datasets.
+In-Distribution (IID) Performance is measured after transfer to the pretraining dataset (ImageNet1k).
+Out-of-Distribution (OOD) Performance is evaluated on datasets not seen during pretraining (iNaturalist18, CIFAR-100, CIFAR-10) (Van Horn et al., 2018), (Krizhevsky, 2009).
+Figure 6: Ensembling Boosts ResNet50 Transfer Accuracy Without Extra Compute Baseline: A single (ResNet50Cat1) model is trained for 400k iters on ImageNet1k (plus 5050k iters for a stronger argument).
+Ensemble: A (ResNet50Cat4) model contains four ResNet50 models trained separately (with different random seeds) on ImageNet for 100k iters each.
+The ensemble significantly outperforms the baseline during transfer.
+22footnotetext: The baseline is pretrained for 450k training steps for a stronger argument.
 5.1.3 Results
-Our method yield considerable improvements in performance despite maintaining the same total pretraining computation.
-By combining more networks that have been pretrained for fewer steps, our experiments produce models which generalize significantly better to new datasets: ($+12.5\%$) on iNaturalist18 and ($+6\%$) on CIFAR-100.
-As expected, adding more sets of features does not significantly improve performance ImageNet1K.
-Since ResNet50 models learn features that provide similar information to reduce training loss, combining them on the training data does not does not yield further improvements.
+Our method yields considerable improvements in performance at fixed training cost.
+On ImageNet1k (IID), the ensemble models performed comparably or slightly worse than the single baseline model, showing no difference in their knowledge of ImageNet1k.
+However, on OOD (Green), the undertrained ensemble significantly outperforms the fully-trained baseline ($12.5\%$ on iNaturalist18, $6\%$ on CIFAR-100).
+This performance gap suggests that an ensemble of undertrained models captures patterns relevant to new tasks which a single fully-trained model may miss.
+In summary, for a fixed computational budget, an ensemble of intentionally under-trained models outperforms a single, fully-trained model when transferred to unseen distributions.
+As expected, concatenating additional sets of features does not significantly boost model performance on ImageNet1K.
+ResNet50 models generate different sets of features which provide similar information to reduce training loss.
+Thus, combining them on ImageNet yields no further improvements.
 If concatenating more models trained on fewer steps can yield a significantly better performing model on new datasets, there may be considerable value left on the table.
-Table 3: Transfer Performance of Concatenated Models, Each Trained For Fewer Iterations A baseline ResNet50 model (cat1) is trained for 450k steps on ImageNet1k.
-Each experimental method (catn) creates a models by concatenating multiple (n) ResNet50 networks, each trained independently for 400k/n steps.
+See Table 3 for detailed results.
+Table 3: Transfer Performance of Concatenated ResNet models with Fixed Total Compute A baseline ResNet50 model (cat1) is trained for 450k iters on ImageNet1k.
+Each experimental method (catn) creates a models by concatenating multiple (n) ResNet50 networks, each trained independently for 400k/n iters.
 Overall computational effort of pretraining is maintained.
-Pretraining Steps Transfer via LinProb
+Pretraining Iters Transfer via LinProb
 Method Per ResNet50 Total ImageNet1k Inat18 Cifar100 Cifar10
@@
 cat5 80k 400k 77.8% 46.2% 77.1% 92.8%
+v
 6 Limitations
@@
   • Group-DRO on water-bird, CelebA, MultiNLI tasks [Sagawa et al., 2020]
+Appendix B Connections to Multi-Calibration
+Missing features fundamentally limits the ability of a model to generalize to new distributions outside the training distribution (OOD).
+[Wald et al., 2021] finds a link between OOD performance and multi-calibration, introduced by [Hebert-Johnson et al., 2018].
+Specifically, they argue that calibration leads to better OOD generalization, by creating invariant representations [Arjovsky et al., 2019].
+[Guy et al., 2025] successfully demonstrate this empirically for in-distribution generalization (IID).
+However, any remediation requires recovering these missing features.